OpenStudy (anonymous):

Please verify Solve log(1-x) = log10 (1-x) = 10 -x = 9 X=9

OpenStudy (phi):

if -x = 9 then x = -9 (not +9) you could solve -x = 9 by multiplying both sides by -1 (or divide both sides by -1, or add -9 to both sides, then +x to both sides)

OpenStudy (anonymous):

Missed typing the negative sign , is -9 correct to make log(1-(-9)= log(10)

OpenStudy (phi):

yes. 1 - -9 is 1+9 = 10